OpenStudy (anonymous):

Please help choose the correct simplification of the expression4/y^-3 . 4y3 4/y^3 y^3/4 This expression is already simplified

OpenStudy (sasogeek):

when you have a negative exponent, it means that, the base of that exponent is on the other side of the fraction with the same exponent but positive. so \(\large x^{-2}\) means \(\huge \frac{1}{x^{2}}\) and \(\huge \frac{1}{x^{-2}}=x^2 \) do you understand this?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

wait i think i get it, so im just going to reveres x^2 into x^-2

OpenStudy (sasogeek):

when you want to reverse x^2, it becomes 1/x^-2

OpenStudy (sasogeek):

it's not called "reverse" anyway but i hope u get it now :)

OpenStudy (anonymous):

i do thank you

OpenStudy (sasogeek):

can you tell what the answer to your question is now?